OpenStudy (anonymous):

what is the domain of 1/ (X-square root of 2) ?

OpenStudy (mr.math):

Is it \(\large \frac{1}{x-\sqrt{2}}?\)

OpenStudy (anonymous):

\[1 \over x - \sqrt{2}\]what value of x makes the denominator equal to zero?

OpenStudy (anonymous):

yes and i have no idea eashmore

OpenStudy (mr.math):

As eashmore said, you have to find the value that makes the denominator zero, that is \[x-\sqrt{2}=0 \implies x=\sqrt{2}.\]

OpenStudy (mr.math):

Now, the domain is real numbers except that point at which the denominator is zero. So the domain is \((-\infty,\infty)-\{\sqrt{2}\}\).

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

After we find what values make the denominator equal to zero, we know that x can equal any other value because\[{1 \over 0}\]is undefined.